# Brain-inspired Large Language Model

### v1

@@ 3 BriLLM Formulation
$\mathcal{E}_{i}$ $=\sum_{k=1}^{i}\mathcal{W}_{k}e_{k}$ (3)
@@
$e_{i+1,v}={\rm GeLU}(W_{x_{i},v}\mathcal{E}_{i}+b_{x_{i},v})$
@@ 4 Experiments
BriLLM is implemented using PyTorch. It uses sinusoidal positional encoding, GeLU as the activation function, cross-entropy loss for next-token prediction, and an embedding size of $d_{model}=32$. We used the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$. The model size is about $512+4000*4000*(32*32+32)\approx 16B$. We trained our models on one machine with 8 NVIDIA A800 GPUs for 1.5k steps. The training loss is shown in Figure 4.
@@ 5 Conclusion, Limitation and the Future
Our existing BriLLM implementation has a size of $d_{model}\times d_{model}\times n\times n$, where $n$ is the number of tokens (nodes) and $d_{model}$ is embedding size. This quadratically increasing model size is indeed an inconvenience. However, as most model parameters come from the fully-connected matrices, it is possible to adopt a sort of sparse representation or shared parameters for those less active tokens, i.e., set a default non-updated matrix for all these inactive tokens. Our preliminary estimation shows such a strategy may save up to 90% parameters for BriLLM.
Both our BriLLM training practice and the SiFu mechanism show BriLLM is hard to efficiently train in parallel as every time the training has to be conducted in a different individual neural network. In addition, a theoretically accurate training objective needs the right predicted token to compare its energy to all the other tokens. When the token set is large, such ranking may result in a very wide softmax output layer, which further slows the training down and requires much larger training memory. It is lucky that such inconvenience may be alleviated by some sort of approximated ranking strategy. Namely, BriLLM training may be done locally only within those ‘necessary’ compared to counterpart tokens. When all these locally trained networks do not overlap, then all these local networks can be trained parallelly, so that the entire BriLLM model training can be done in a good parallel way.
@@
Note that even though BriLLM theoretically supports an infinite-gram language model without increasing model size, in practice, the model during training has to cover long enough input sequences, otherwise BriLLM decoding cannot give good enough sequence prediction beyond the training sample length. However, facilitating longer sequence prediction in terms of BriLLM just depends on longer training without resizing the model itself.

### v2

@@ 3 BriLLM Formulation
$\mathcal{E}_{i}$ $=\sum_{k=1}^{i}\mathcal{W}_{k}e_{k},$ (3)
where $\mathcal{W}$ represents the softmax-normalized weights.
@@
$e_{i+1,v}={\rm GeLU}(W_{x_{i},v}\mathcal{E}_{i}+b_{x_{i},v}+PE_{i}),$
where $PE$ represents the sine and cosine positional encoding.
@@ 4 Experiments
BriLLM is implemented using PyTorch. It uses sine and cosine positional encoding, GeLU as the activation function, cross-entropy loss for next-token prediction, and an embedding size of $d_{model}=32$. We used the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$. The model size is about $512+4000*4000*(32*32+32)\approx 16B$. We trained our models on one machine with 8 NVIDIA A800 GPUs for 1.5k steps. The training loss is shown in Figure 4.
@@ 5 Conclusion, Limitation and the Future
Our existing BriLLM implementation has a size of $d_{model}\times d_{model}\times n\times n$, where $n$ is the number of tokens (nodes) and $d_{model}$ is embedding size. This quadratically increasing model size is indeed inconvenience. However, as most model parameters come from the fully-connected matrices, it is possible to adopt a sort of sparse representation or shared parameters for those less active tokens, i.e., set a default non-updated matrix for all these inactive tokens. Our preliminary estimation show such strategy may save up to 90% parameters for BriLLM.
Both our BriLLM training practice and the SiFu mechanism show BriLLM is hard to efficiently trained in parallel as every time the training has to be conducted in a different individual neural network. In addition, theoretically accurate training objective needs the right predicted token has to compared its energy to all the other tokens. When token set is large, such ranking may result in a very wide softmax output layer, which further slows the training down and requires much larger training memory. It is lucky that such inconvenience may be alleviated by some sort of approximated ranking strategy. Namely, BriLLM training may be done locally only within those ‘necessary’ compared counterpart tokens. When all these locally trained networks does not overlap, then all these local network can be trained parallelly, so that the entire BriLLM model training can be done in a good parallel way.
@@
Note that even though BriLLM theoretically supports infinite-gram language model without increasing model size, in practice, the model during training has to cover long enough input sequences, otherwise BriLLM decoding cannot give good enough sequence prediction beyond the training sample length. However, facilitating longer sequence prediction in terms of BriLLM just depends on longer training without resizing the model itself.